If ΔG∘ for a reaction equals −5.92 * 105 J/mol, what is K? Assume T =...

If ΔG for a reaction equals −5.92 * 105 J/mol, what is K? Assume T = 25
Answer #1

T= 25.0 oC

= (25.0+273) K

= 298 K

ΔGo = -592000 J/mol

use:

ΔGo = -R*T*ln Kc

-592000 = - 8.314*298.0* ln(Kc)

ln Kc = 238.9436

Kc = 5.914*10^103

Answer: 5.91*10^103